# ClavaDDPM: Multi-relational Data Synthesis with Cluster-guided Diffusion Models

### v1

@@ Abstract
Recent research in tabular data synthesis has focused on single tables, whereas real-world applications often involve complex data with tens or hundreds of interconnected tables. Previous approaches to synthesizing multi-relational (multi-table) data fall short in two key aspects: scalability for larger datasets and capturing long-range dependencies, such as correlations between attributes spread across different tables. Inspired by the success of diffusion models in tabular data modeling, we introduce Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM). This novel approach leverages clustering labels as intermediaries to model relationships between tables, specifically focusing on foreign key constraints. ClavaDDPM leverages the robust generation capabilities of diffusion models while incorporating efficient algorithms to propagate the learned latent variables across tables. This enables ClavaDDPM to capture long-range dependencies effectively. Extensive evaluations on multi-table datasets of varying sizes show that ClavaDDPM significantly outperforms existing methods for these long-range dependencies while remaining competitive on utility metrics for single-table data.
@@ 1 Introduction
In this work, we: 1) provide a complete formulation of the multi-relational modeling process, as well as the essential underlying assumptions being made, 2) propose an efficient framework to generate multi-relational data that preserves long-range dependencies between tables, 3) propose relation-aware clustering as a proxy for modeling parent-child constraints, and apply the controlled generation capabilities of diffusion models to tabular data synthesis, 4) apply an approximate nearest neighbor search-based matching technique, as a universal solution to the multi-parent relational synthesis problem for a child table with multiple parents, 5) establish a comprehensive multi-relational benchmark, and propose long-range dependency as a new metric to measure synthetic data quality specific to multi-table cases, and 6) show that ClavaDDPM significantly outperforms existing methods for these long-range dependency metrics while remaining competitive on utility metrics for single-table data.
@@ 3 Background
A multi-relational database $\mathcal{R}$ consists of $m$ tables $(R_{1},\ldots,R_{m})$. Each table is a collection of rows, which are defined over a sequence of attributes. One of the attributes, let’s consider the first attribute without loss of generality, is the primary key of table $R$, which serves as the unique identifier for each row in the table. No rows in the same table have repeated values for the primary key attribute. We use Berka database [4] as our running example in this work, as in Figure 1. Note the Account ID, the primary key for the Account table in Berka.
Given a table $R_{j}$, we say a relation $R_{i}$ has a foreign key constraint with $R_{j}$ or $R_{i}$ refers to $R_{j}$, if $R_{i}$ has an attribute known as foreign key that refers to the primary key of $R_{j}$: for every row $r_{i}\in R_{i}$, there exists a row $r_{j}\in R_{j}$ such that $r_{j}$’s primary key value equals to $r_{i}$’s foreign key value. For example, the Account ID of the Loan table refers to the primary key of the Account table. If an account row is removed from the Account ID table, so would all the referring rows in the Loan table to this account, for foreign key constraint to hold. Note that the primary key of a table can consist of multiple attributes. In this paper, we focus on the case of a single attribute that is common in practice.
@@
$q\left(\boldsymbol{x}_{t}\left.\right|\boldsymbol{x}_{t-1}\right)\coloneq\mathcal{N}\left(\boldsymbol{x}_{t};\sqrt{1-\beta_{t}}\boldsymbol{x}_{t-1},\beta_{t}\boldsymbol{I}\right),p_{\theta}\left(\boldsymbol{x}_{t-1}\left.\right|\boldsymbol{x}_{t}\right)\coloneq\mathcal{N}\left(\boldsymbol{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\boldsymbol{x}_{t},t\right),\boldsymbol{\Sigma}_{\theta}\left(\boldsymbol{x}_{t},t\right)\right).$ (2)
@@
By approximating $\log p_{\phi}\left(\boldsymbol{y}\left.\right|\boldsymbol{x}_{t}\right)$ using Taylor expansion around $\boldsymbol{x}_{t}=\boldsymbol{\mu}$, the conditional reverse process (3) can be approximated with a perturbed Gaussian transition [10]
@@ 4.1 Modeling generative process for two-table relational databases
Consider a database of two tables =$\mathcal{R}=\{R_{1},R_{2}\}$, e.g. {Loan, Account} in Berka, where the child table $R_{1}$ refers to parent table $R_{2}$. To model the entire database, we first use $\boldsymbol{X}$ and $\boldsymbol{Y}$ as the variables for the child table $R_{1}$ and parent table $R_{2}$, respectively (dropping their primary key attributes and indexing their respective row variables starting from one). In this section, we use boldface to represent random variables. e.g. $Y\sim\boldsymbol{Y}$, where $Y$ is the data of $R_{2}$, and $\boldsymbol{Y}$ is the random variable $Y$ being sampled from. In addition, we use subscript to represent the parent row some data or random variable refers to. e.g. $\boldsymbol{x}_{j}$ represents the child random variable who refers to parent $\boldsymbol{y}_{j}$.
@@
2) The i.i.d assumption does not apply to the child table rows ($\boldsymbol{x}_{j}$’s) as they are constrained by their respective parent rows. Consider two loans associated with the same account id; if one’s status is in debt (“C”), the other one is likely so too. However, we can assume that the child rows referring to different parent rows are independent. Hence, we model $\boldsymbol{X}$ by $\left\{\boldsymbol{g}_{j}\left.\right|j=1,\ldots,\left|R_{2}\right|\right\}$, where each group $\boldsymbol{g}_{j}=\left\{\boldsymbol{x}_{j}^{i}\left.\right|i=1,\ldots,\left|\boldsymbol{g}_{j}\right|\right\}$ represents a set of child table rows referring to the parent row $\boldsymbol{y}_{j}$.
@@
where $X=\cup_{j=1}^{|R_{2}|}g_{j}$ and $g_{j}=\{x_{j}^{1},\ldots,x_{j}^{|g_{j}|}\}$. This model allows us to capture the inter-table correlations (the correlation between tuples between different tables) and the intra-group correlations.
@@
Phase I: Latent learning and table augmentation: (1) Learn latent variable $c$ on the joint space $\left(X;Y\right)$, such that each parent row $y_{j}$ corresponds to a learned latent variable $c_{j}$. (2) Augment the parent table into $T_{Y}=\left(Y;C\right)$, where $C$ corresponds the latent variable values $c_{j}$ for each row $y_{j}$.
@@ 4.3 Design choices for ClavaDDPM
4.3.1 Relation-aware clustering
@@
However, such clustering in the joint space $\left(\boldsymbol{X};\lambda\boldsymbol{Y}\right)$ could potentially lead to inconsistency when we create the augmented table $T_{Y}=\left(Y;C\right)$. Though we add a weight $\lambda$ to the parent rows such that child rows with the same parent rows are likely to be assigned to the same cluster, there is still some chance that they end with different clusters. In particular, for each parent row $y_{j}\in Y$, its child rows are assigned to different clusters. In LavaDDPM, we impose a majority voting step to find the most popular cluster label in each foreign key group and assign it to the parent row $y_{j}$. In practice, the voting agree rates tend to be high, and this can be further enforced by assigning a higher weight to the parent table (increasing $\lambda$) during GMM clustering. We evaluate the choice of $\lambda$ and voting agree rates in our ablation study in Section 5.3.
@@ 5.2 End-to-end evaluation
We conducted multi-table synthesis experiments on five multi-table datasets and report the averaged utility with standard deviation for all algorithms in Table 1. First, the evaluation shows that ClavaDDPM has an overall advantage against all the baseline models in terms of correlation modeling, and is surpassing the baselines by larger margins for longer-range dependencies. e.g. in Instacart 05, our model outperforms the best baseline by $58.29\%$ on $2$-hop correlations, and in Berka, our model exceeds the best baseline by $20.24\%$ on $3$-hop correlations. For single-column densities and cardinality distributions, ClavaDDPM exhibits a competitive result compared to the state-of-the-art baseline models. We also evaluate ClavaDDPM against baselines on high-order single-table metrics (Appendix 5.3), which shows that our model has advantages in preserving data fidelity, generating diverse data, and achieving high machine learning efficacy.
@@
Table 1: End-to-end results. DNC denotes Did Not Converge, and TLE denotes Time Limit Exceeded. ST stands for Single-T and D stands for Dnorm.
@@ 5.3 Ablation study
avg agree-rate $43.99$
± $\,\pm\,0.81$
$100.00$
± $\,\pm\,0.00$
$28.84$
± $\,\pm\,0.14$
$44.97$
± $\,\pm\,4.02$
$43.74$
± $\,\pm\,0.60$
$45.32$
± $\,\pm\,0.63$
NA NA
Table 2: Ablation studies on number of clusters $k$, parent scale $\lambda$, and classifier gradient scale $\eta$.
Number of clusters $k$. We study the necessity of using latent cluster conditioning: (i) no conditioning with $k=1$; (ii) many clusters with $k=1000$ to approximate a direct conditioning on parent rows rather than latent variables. When $k=1$, the quality of long-range correlation degrades drastically. When $k=1000$, we still get reasonably strong performance, which showcases ClavaDDPM’s robustness. Compared to the default setting ($k=20$), the metrics are lower in all of cardinality distribution, single column densities, and column correlations — proper latent variable learning leads to better results than direct conditioning on parent rows. We also report a new metric avg agree-rate, the average of all per-table agree rates for the labels within each foreign key group (Section 4.3.1). This measure highly depends on $k$, but a higher rate does not always imply a better performance (e.g., k=1 achieves perfect rates). We provide more insights on how it varies with the next parameter.
Parent scale $\lambda$. Varying the parent scale parameter $\lambda$ changes the agree-rates as shown in Table 2, but the downstream model performance do not vary too much. This result indicates that the relation-aware clustering process is robust against such factors, and the GMM model is capable of capturing nuances in data patterns. The detailed discussion is in Appendix D.1.
@@ 6 Conclusion
We focused on foreign key constraint in this work, and assumed that the parent-child relationship is known among the tables as well as the common assumption that parent rows or the child rows corresponding to different primary keys are i.i.d. This brings three natural follow-up research directions: i) extension to the scenarios where this prior information is not available and these relationships need to be discovered first [29], ii) relaxing the i.i.d. assumptions, and iii) inspecting multi-relational data synthesis with other integrity constraints (e.g, denial constraints [15], general assertions for business rules).
Furthermore, we evaluated ClavaDDPM’s privacy with the common (in tabualr data literature) DCR metric. Nonetheless, we think it is worthwhile to: i) evaluate the resiliency of ClavaDDPM against stronger privacy attacks [43] , and ii) investigate the efficacy of boosting ClavaDDPM with privacy guarantees such as differential privacy. Similarly, the impacts of our design on fairness and bias removal, as another motivating pillar in synthetic data generation, is well worth exploring as future work. We believe the through multi-relational modeling formulation we presented in this work, can serve as a strong foundation to build private and fair solutions upon.
@@ Appendix B Algorithm Details
B.1 Multi-relational synthesis problem
Given a multi-relational database =$\mathcal{R}=\{R_{1},\ldots,R_{m}\}$, we would like to generate a synthetic version ~=$\tilde{\mathcal{R}}=\{\tilde{R}_{1},\ldots,\tilde{R}_{m}$} that has the same structure and foreign-key constraints as $\mathcal{R}$ and preserves attribute correlations within $\mathcal{R}$, including 1) the inter-column correlations within the same table; 2) the intra-group correlations within the same foreign key group; 3) the inter-table correlations. The first aspect has been well defined, measured, and tackled in the literature of single-table synthesis [52, 25, 50] while the other two aspects are raised due to foreign-key constraints between tables [5].
B.2 Diagram for two-table relational databases

### v2

@@ Abstract
Recent research in tabular data synthesis has focused on single tables, whereas real-world applications often involve complex data with tens or hundreds of interconnected tables. Previous approaches to synthesizing multi-relational (multi-table) [^1] In the context of databases, a table is formally referred to as a relation. Throughout this work, we use these terms interchangeably. data fall short in two key aspects: scalability for larger datasets and capturing long-range dependencies, such as correlations between attributes spread across different tables. Inspired by the success of diffusion models in tabular data modeling, we introduce Cluster Latent Variable guided Denoising Diffusion Probabilistic Models (ClavaDDPM). This novel approach leverages clustering labels as intermediaries to model relationships between tables, specifically focusing on foreign key constraints. ClavaDDPM leverages the robust generation capabilities of diffusion models while incorporating efficient algorithms to propagate the learned latent variables across tables. This enables ClavaDDPM to capture long-range dependencies effectively. Extensive evaluations on multi-table datasets of varying sizes show that ClavaDDPM significantly outperforms existing methods for these long-range dependencies while remaining competitive on utility metrics for single-table data.
@@ 1 Introduction
In this work, we: 1) provide a complete formulation of the multi-relational modeling process, as well as the essential underlying assumptions being made, 2) propose an efficient framework to generate multi-relational data that preserves long-range dependencies between tables, 3) propose relationship-aware clustering as a proxy for modeling parent-child constraints, and apply the controlled generation capabilities of diffusion models to tabular data synthesis, 4) apply an approximate nearest neighbor search-based matching technique, as a universal solution to the multi-parent relational synthesis problem for a child table with multiple parents, 5) establish a comprehensive multi-relational benchmark, and propose long-range dependency as a new metric to measure synthetic data quality specific to multi-table cases, and 6) show that ClavaDDPM significantly outperforms existing methods for these long-range dependency metrics while remaining competitive on utility metrics for single-table data.
@@ 3 Background
A multi-relational database $\mathcal{R}$ consists of $m$ tables (or relations) $(R_{1},\ldots,R_{m})$. Each table is a collection of rows, which are defined over a sequence of attributes. One of the attributes, let’s consider the first attribute without loss of generality, is the primary key of table $R$, which serves as the unique identifier for each row in the table. No rows in the same table have repeated values for the primary key attribute. We use Berka database [4] as our running example in this work, as in Figure 1. Note the Account ID, the primary key for the Account table in Berka.
Given a table $R_{j}$, we say a relation $R_{i}$ has a foreign key constraint with $R_{j}$, or $R_{i}$ refers to $R_{j}$, if $R_{i}$ has an attribute known as foreign key that refers to the primary key of $R_{j}$: for every row $r_{i}\in R_{i}$, there exists a row $r_{j}\in R_{j}$ such that $r_{j}$’s primary key value equals to $r_{i}$’s foreign key value. For example, the Account ID of the Loan table refers to the primary key of the Account table. If an account row is removed from the Account ID table, so would all the referring rows in the Loan table to this account, for foreign key constraint to hold. Note that the primary key of a table can consist of multiple attributes. In this paper, we focus on the case of a single attribute that is common in practice. Also note that all keys are considered row identifiers and are thus not treated or modeled alongside the actual table attributes in this work.
@@
Multi-relational synthesis problem.
Given a multi-relational database =$\mathcal{R}=\{R_{1},\ldots,R_{m}\}$, we would like to generate a synthetic version ~=$\tilde{\mathcal{R}}=\{\tilde{R}_{1},\ldots,\tilde{R}_{m}$} that has the same structure and foreign-key constraints as $\mathcal{R}$ and preserves attribute correlations within $\mathcal{R}$, including 1) the inter-column correlations within the same table; 2) the intra-group correlations within the same foreign key group; 3) the inter-table correlations. The first aspect has been well defined, measured, and tackled in the literature of single-table synthesis [52, 25, 50] while the other two aspects are raised due to foreign-key constraints between tables [5]. For instance, in Berka database (Figure 1), the foreign key constraint between the Loan table and the Account table via Account ID adds an important intra-group correlation for the combinations of loans associated with an account and many 1-hop inter-table correlations between columns in the Loan table and the columns in the Account table. Even for the Loan table and the Demographic table that are indirectly constrained by foreign keys, their columns are correlated as well, e.g., how is the average salary in a district related to the status of loans, an example for 2-hop inter-table correlation.
@@
$q\left(\boldsymbol{x}_{t}\left.\right|\boldsymbol{x}_{t-1}\right)$ :-$\coloneq\mathcal{N}\left(\boldsymbol{x}_{t};\sqrt{1-\beta_{t}}\boldsymbol{x}_{t-1},\beta_{t}\boldsymbol{I}\right):$ (2)
$p_{\theta}\left(\boldsymbol{x}_{t-1}\left.\right|\boldsymbol{x}_{t}\right)$ :-$\coloneq\mathcal{N}\left(\boldsymbol{x}_{t-1};\boldsymbol{\mu}_{\theta}\left(\boldsymbol{x}_{t},t\right),\boldsymbol{\Sigma}_{\theta}\left(\boldsymbol{x}_{t},t\right)\right).:$
@@
By approximating $\log p_{\phi}\left(\boldsymbol{y}\left.\right|\boldsymbol{x}_{t}\right)$ using Taylor expansion around $\boldsymbol{x}_{t}=\boldsymbol{\mu}$, the conditional reverse process (Equation 3) can be approximated with a perturbed Gaussian transition [10]
@@ 4.1 Modeling generative process for two-table relational databases
Consider a database of two tables =$\mathcal{R}=\{R_{1},R_{2}\}$, e.g. {Loan, Account} in Berka, where the child table $R_{1}$ refers to parent table $R_{2}$. To model the entire database, we first use $\boldsymbol{X}$ and $\boldsymbol{Y}$ as the variables for the child table $R_{1}$ and parent table $R_{2}$, respectively (dropping their primary key attributes and indexing their respective row variables starting from one). In this section, we use boldface to represent random variables. e.g. $Y\sim\boldsymbol{Y}$, where $Y$ is the data of $R_{2}$, and $\boldsymbol{Y}$ is the random variable $Y$ being sampled from. In addition, we use subscript to represent the parent row some data or random variable refers to. e.g. $\boldsymbol{x}_{j}$ represents the child random variable who refers to parent $\boldsymbol{y}_{j}$. Refer to Appendix A for a complete list of notations used and the corresponding design choices.
@@
2) The i.i.d assumption does not apply to the child table rows ($\boldsymbol{x}_{j}$’s) as they are constrained by their respective parent rows. Consider two loans associated with the same account id; if one’s status is in debt (“C”), the other one is likely so too. To capture this dependency, we make a Bayesian modeling assumption that, although child rows associated with the same parent row are not independent, they are conditionally independent of child rows associated with other parent rows, given their respective parent. For example, consider an account table (parent) and a loan table (child). Loans related to the same account (i.e., the same parent) are not independent due to shared account-specific factors. However, loans from different accounts can be considered conditionally independent when accounting for their respective account-level information. Hence, we model $\boldsymbol{X}$ by $\left\{\boldsymbol{g}_{j}\left.\right|j=1,\ldots,\left|R_{2}\right|\right\}$, where each group $\boldsymbol{g}_{j}=\left\{\boldsymbol{x}_{j}^{i}\left.\right|i=1,\ldots,\left|\boldsymbol{g}_{j}\right|\right\}$ represents a set of child table rows referring to the parent row $\boldsymbol{y}_{j}$.
@@
where $X=\cup_{j=1}^{|R_{2}|}g_{j}$ and $g_{j}=\{x_{j}^{1},\ldots,x_{j}^{|g_{j}|}\}$. This model allows us to capture the inter-table correlations (the correlation between tuples from different tables) and the intra-group correlations.
@@
Phase I: Latent learning and table augmentation: (1) Learn latent variable $c$ on the joint space $\left(X;Y\right)$, such that each parent row $y_{j}$ corresponds to a learned latent variable $c_{j}$. (2) Augment the parent table into $T_{Y}=\left(Y;C\right)$, where $C$ corresponds to the latent variable values $c_{j}$ for each row $y_{j}$.
@@ 4.2 Extension to more parent-child constraints
Extending the model to include more tables allows for capturing longer-range dependencies, beyond just those between adjacent tables. For example, as shown in Figure 1, the dependency between the Demographic table and the Credit Card table can also be captured and quantified. Further details are provided in Section 5.
@@ 4.3 Design choices for ClavaDDPM
4.3.1 Relationship-aware clustering
@@
However, such clustering in the joint space $\left(\boldsymbol{X};\lambda\boldsymbol{Y}\right)$ could potentially lead to inconsistency when we create the augmented table $T_{Y}=\left(Y;C\right)$. Though we add a weight $\lambda$ to the parent rows such that child rows with the same parent rows are likely to be assigned to the same cluster, there is still some chance that they end with different clusters. In particular, for each parent row $y_{j}\in Y$, its child rows are assigned to different clusters. In ClavaDDPM, we impose a majority voting step to find the most popular cluster label in each foreign key group and assign it to the parent row $y_{j}$. In practice, the voting agree rates tend to be high, and this can be further enforced by assigning a higher weight to the parent table (increasing $\lambda$) during GMM clustering. We evaluate the choice of $\lambda$ and voting agree rates in our ablation study in Section 5.3.
@@ 5.2 End-to-end evaluation
We conducted multi-table synthesis experiments on five multi-table datasets and report the averaged utility with standard deviation for all algorithms in Table 1. First, the evaluation shows that ClavaDDPM has an overall advantage against all the baseline models in terms of correlation modeling, and is surpassing the baselines by larger margins for longer-range dependencies. e.g. in Instacart 05, our model outperforms the best baseline by $58.29\%$ on $2$-hop correlations, and in Berka, our model exceeds the best baseline by $20.24\%$ on $3$-hop correlations. For single-column densities and cardinality distributions, ClavaDDPM exhibits a competitive result compared to the state-of-the-art baseline models. We also evaluate ClavaDDPM against baselines on high-order single-table metrics (Appendix D.3), which shows that our model has advantages in preserving data fidelity, generating diverse data, and achieving high machine learning efficacy.
@@
Table 1: End-to-end results. DNC denotes Did Not Converge, and TLE denotes Time Limit Exceeded. ST stands for Single-T and D stands for Denorm. Statistical metrics described in Section 5.1 are reported.
@@ 5.3 Ablation study
avg agree-rate $81.12$
± $\,\pm\,0.99$
$100.00$ $\,\pm\,0.00$ $65.97$
± $\,\pm\,0.17$
$78.85$
± $\,\pm\,0.73$
$80.87$
± $\,\pm\,0.73$
$82.44$
± $\,\pm\,0.77$
$81.32$
± $\,\pm\,0.80$
$81.37$
± $\,\pm\,1.09$
$80.88$
± $\,\pm\,0.58$
Table 2: Ablation studies on number of clusters $k$, parent scale $\lambda$, and classifier gradient scale $\eta$. Note that $\eta$ and matching have no effect on agree rates. Statistical metrics described in Section 5.1 are reported.
Number of clusters $k$. We study the necessity of using latent cluster conditioning: (i) no conditioning with $k=1$; (ii) many clusters with $k=1000$ to approximate a direct conditioning on parent rows rather than latent variables. When $k=1$, the quality of long-range correlation degrades drastically. When $k=1000$, we still get reasonably strong performance, which showcases ClavaDDPM’s robustness. Compared to the default setting ($k=20$), the metrics are lower in all of cardinality distribution, single column densities, and column correlations — proper latent variable learning leads to better results than direct conditioning on parent rows. We also report a new metric avg agree-rate, the average of all per-table agree rates for the labels within each foreign key group (Section 4.3.1). This measure highly depends on $k$, but a higher rate does not always imply a better performance (e.g., k=1 achieves perfect rates). We provide more insights on how it varies with the next parameter. We also conducted finer-grained experiments to examine the effect of $k$ on model performance, as shown in Appendix D.2.
Parent scale $\lambda$. Varying the parent scale parameter $\lambda$ changes the agree-rates as shown in Table 2, but the downstream model performance does not vary too much. This result indicates that the relation-aware clustering process is robust against such factors, and the GMM model is capable of capturing nuances in data patterns. The detailed discussion is in Appendix D.1.
@@ 6 Conclusion
We focused on foreign key constraints in this work, and made the assumption that child rows are conditionally independent given corresponding parent rows. This brings three natural follow-up research directions: i) extension to the scenarios where this prior information is not available and these relationships need to be discovered first[29], ii) further relaxing the assumptions, and iii) inspecting multi-relational data synthesis with other integrity constraints (e.g, denial constraints[15], general assertions for business rules). Furthermore, we evaluated ClavaDDPM’s privacy with the common (in tabular data literature) DCR metric. Nonetheless, we think it is worthwhile to: i) evaluate the resiliency of ClavaDDPM against stronger privacy attacks[43], and ii) investigate the efficacy of boosting ClavaDDPM with privacy guarantees such as differential privacy. Similarly, the impacts of our design on fairness and bias removal, as another motivating pillar in synthetic data generation, is well worth exploring as future work. We believe the thorough multi-relational modeling formulation we presented in this work, can serve as a strong foundation to build private and fair solutions upon.
Acknowledgments
This work was supported by NSERC through a Discovery Grant, an alliance grant, the Canada CIFAR AI Chairs program. Resources used in preparing this research were provided, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute. We thank the reviewers and program chairs for their detailed comments, which greatly improved our paper.
@@ Appendix B Algorithm Details
B.1 Diagram for two-table relational databases